16. Evaluate the following: [2x(x²-9°dx
Answer #1

16)To evalaute the definite integral , \int_{0}^{2}2x(x^2-9)^3dx , expand the integrand as

\int_{0}^{2}2x(x^2-9)^3dx=2\int_{0}^{2}x(x^6-27x^4+243x^2-729)dx\\ \int_{0}^{2}2x(x^2-9)^3dx=2\int_{0}^{2}(x^7-27x^5+243x^3-729x)dx\\ \int_{0}^{2}2x(x^2-9)^3dx=2(x^8/8-27x^6/6+243x^4/4-729x^2/2]_{0}^{2}\\ {\color{Blue} \int_{0}^{2}2x(x^2-9)^3dx=-1484}